OpenStudy (anonymous):

Evaluate (ab)2 for a = 4 and b = 3.

OpenStudy (anonymous):

Is that "(ab) times 2" or \[(ab)^{2}\]

OpenStudy (baldymcgee6):

simply replace a with 4 and b with 3 in the given expression

OpenStudy (anonymous):

isnt that just like 86

OpenStudy (baldymcgee6):

given that (ab)2 where a =4 and b= 3 (4*3)2= (12)2 = 24

OpenStudy (anonymous):

still wooda done it my way

OpenStudy (anonymous):

yeah thanks yall

OpenStudy (baldymcgee6):

if your way gave you 86, then it is wrong.

OpenStudy (anonymous):

if it is really (ab)^2, then both 24 and 86 are wrong... ( (4)(3) )^2 = 12^2 = 144

OpenStudy (baldymcgee6):

Assuming he posted the original question correctly.